OpenStudy (anonymous):

3/7(14 − 7c) − 9 ≥3/5(10c − 15) + 18

OpenStudy (anonymous):

expand first

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i was trying to figure that one out for them

OpenStudy (anonymous):

6-3c-9>=6c-9+18 -3c-3>=6c+9 -12>=9c (-12/9)>=c (-4/3)>=c c<=(-4/3)